A study reported that in a random sampling of 100 women over the age of 35 showed that 8 of the women were married 2 or more times. In a group of 5,000 women based on the study results, how many would likely be married 2 or more times?
Answer by stanbon(48568)

A study reported that in a random sampling of 100 women over the age of 35 showed that 8 of the women were married 2 or more times. In a group of 5,000 women based on the study results, how many would likely be married 2 or more times?
--------
It looks like a proportion problem and not a statistics problem:
x/5000 = 8/100
x/50=8/1
x=400
Cheers,
Stan H.